OpenStudy (anonymous):

Let z be a complex number than locus represented by |iz-1| + |z-i| = 2 is : a) a line , b) a circle , c) a pair of straight lines, d) a parabola

OpenStudy (anonymous):

i think it is a l ine

OpenStudy (anonymous):

Sorry for the mistake in the question earlier , the modified quest. is in the post itself. I seek help from the users presented here as soon as possible,

OpenStudy (anonymous):

@satellite73 rethink please :)

OpenStudy (anonymous):

scratch that, i think it is a region bounded by two line

OpenStudy (anonymous):

OK so since it is IIT based so I will say : There may be no answers or multiple answers

OpenStudy (anonymous):

Well yes @satellite73 has the answer, it will be a line \(\textbf{segment}\)

OpenStudy (anonymous):

But I am half way stuck @satellite73 , can you show your work, I think @experimentX is also writing his work :)

OpenStudy (experimentx):

put z = x + iy you would get, \[ \sqrt{(y-1)^2 + x^2} + \sqrt{x^2 + (y-1)^2} = 2\] http://www.wolframalpha.com/input/?i=plot+ \sqrt{%28y-1%29^2+%2B+x^2}+%2B+\sqrt{x^2+%2B+%28y-1%29^2}+%3D+2 guess it is a circle

OpenStudy (experimentx):

woops!! (y+1)^2

OpenStudy (anonymous):

What I can do is : \[\large{|(-iz-1)| + |(z-i)| =2}\] \[\large{|i(z+i)|+|z-i|=2}\] \[\large{|i||z+i| + |z-i| = 2 }\] \[\large{|z+i| + |z-i| = 2}\]

OpenStudy (anonymous):

that looks good

OpenStudy (anonymous):